OpenStudy (anonymous):

Calculate the average rate of change for the given function, from x = 8 to x = 5. x f(x) 8 0 5 4 −5 19 negative four thirds four thirds negative three fourths three fourths Can you please help me out with this one? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

when x = 8, f(x) = ???

OpenStudy (anonymous):

0

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

and when x = 5, f(x) = ???

OpenStudy (anonymous):

4

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

so \[\Large \frac{f(8) - f(5)}{8 - 5} = \frac{0-4}{8-5}\] \[\Large \frac{f(8) - f(5)}{8 - 5} = \frac{-4}{3}\] \[\Large \frac{f(8) - f(5)}{8 - 5} = -\frac{4}{3}\] is the average rate of change

jimthompson5910 (jim_thompson5910):

in other words, you divide the difference in y (f(8) and f(5) with the difference in x (8 and 5) you subtract in the same order to make sure you have the correct sign